In a factory there are three
independently operating machines, each with a
probability of 0.6 of being working, find the probability that:
a. All the machines are working
b. None of the machines are working
c. Only one machine is working

Answer by rothauserc(4718)

a. Probability(P) (all three are working) = 0.6^3 = 0.216
:
b: P(no machines are working) = 0.4^3 = 0.064
:
c: P( only one machine is working) = 3 * 0.60 * 0.40^2 = 0.288
: